OpenStudy (anonymous):

what are the slope and y-intercept of the equation 3x-4y=4 ? the choices are a. slope:-3, y-intercept:-4 b. slope 3/4, y-intercept:-1 c. slope:3, y-intercept:4 d. slope:3/4, y-intercept:-4

hartnn (hartnn):

for \(y= mx+c\) the slope = m and y -intercept = c so first try to isolate y from 3x-4y=4

OpenStudy (anonymous):

I got -y=8-3x

hartnn (hartnn):

how ?

hartnn (hartnn):

3x-4y=4 start by adding 4y on both sides...

OpenStudy (anonymous):

ok done I got 3x=8y

hartnn (hartnn):

you cannot combine 4y and 4 they are different terms \(3x = 4 + 4y\) now subtract 4 from both sides

OpenStudy (anonymous):

ok 3x-4=4y

hartnn (hartnn):

correct1 now divide both sides by 4 :)

OpenStudy (anonymous):

ok

hartnn (hartnn):

what did u get ?

OpenStudy (anonymous):

3/4x=4+y

OpenStudy (anonymous):

I have to get y by it self ?

hartnn (hartnn):

good try, but see this..  did u get that ^ ?